# SeTAR: Out-of-Distribution Detection with Selective Low-Rank Approximation

### v1

@@ Abstract
Out-of-distribution (OOD) detection is crucial for the safe deployment of neural networks. Existing CLIP-based approaches perform OOD detection by devising novel scoring functions or sophisticated fine-tuning methods. In this work, we propose SeTAR, a novel, training-free OOD detection method that leverages selective low-rank approximation of weight matrices in vision-language and vision-only models. SeTAR enhances OOD detection via post-hoc modification of the model’s weight matrices using a simple greedy search algorithm. Based on SeTAR, we further propose SeTAR+FT, a fine-tuning extension optimizing model performance for OOD detection tasks. Extensive evaluations on ImageNet1K and Pascal-VOC benchmarks show SeTAR’s superior performance, reducing the false positive rate by up to 18.95% and 36.80% compared to zero-shot and fine-tuning baselines. Ablation studies further validate our approach’s effectiveness, robustness, and generalizability across different model backbones.
@@ 1 Introduction
Previous research has primarily focused on detecting OOD instances in either visual (DeVries & Taylor, 2018; Hendrycks & Gimpel, 2017; Liang et al., 2018; Hendrycks et al., 2022) or textual data (Hu & Khan, 2021; Zheng et al., 2020; Zhou et al., 2021; Tan et al., 2019). Recently, significant progress has been made in multimodal tasks like multimodal retrieval (Li et al., 2023; Caesar et al., 2018) and image classification (Yu et al., 2022), thanks to vision-and-language pretrained (VLP) models like CLIP (Radford et al., 2021). More recent studies have explored OOD detection with CLIP, grouped into zero-shot methods (Fort et al., 2021; Ming et al., 2022; Miyai et al., 2023b) and finetuning-based methods (Ming & Li, 2023; Tao et al., 2023; Miyai et al., 2023a). However, the zero-shot methods suffer from suboptimal performance due to potential domain gaps with ID downstream data. On the other hand, finetuning-based methods carry the risk of deconstructing the intricate representations learned by CLIP which requires a meticulously designed training strategy.
@@
We conduct extensive evaluations and achieve state-of-the-art performance on common OOD detection benchmarks for CLIP, including the ImageNet1K and Pascal-VOC benchmarks. Compared to vanilla MCM and GL-MCM, SeTAR with the CLIP backbone reduces FPR95 by 9.5% and 12.0% on average across two benchmarks, respectively. When further integrate fine-tuning into SeTAR, SeTAR+FT outperforms the state-of-the-art fine-tuning baselines LoCoOp (Miyai et al., 2023a) and LoRA (Hu et al., 2022). Moreover, we perform a comprehensive ablation study and analysis to verify and understand SeTAR. In summary, our key results and contributions:
@@
    We proposed SeTAR, a simple yet effective OOD detection method based on selective low-rank approximation. It is training-free as it only performs post-hoc modification to CLIP weight matrices. SeTAR applies to a variety of scoring functions and model backbones. It can be readily integrated with existing zero-shot OOD detection methods.
@@ 2 Preliminaries
The CLIP model (Radford et al., 2021) contains an image encoder $E^{v}(\cdot)$ and a text encoder $E^{t}(\cdot)$, aligned by contrastive learning over web-scale image-text pairs. In this work, we focus on CLIP-ViT which has a vision Transformer (ViT) as the image encoder. Each ViT encoder layer consists of a multihead self-attention sublayer and a feed-forward sublayer (FFN). In the self-attention module, the hidden state is linearly projected into different spaces with learnable parameters $\mathrm{W_{q},W_{k},W_{v}}$. The output of the multihead self-attention is concatenated and projected back with another trainable linear matrix $\mathrm{W_{o}}$. In the feed-forward module, the hidden state is projected into a wider space with up-projection matrix $\mathrm{W_{\text{up}}}$ and then projected back using $\mathrm{W_{\text{down}}}$ after the non-linear activation function (Figure 1). Due to the similarity of the image and text encoder layers, we use the same notations for the linear matrices in these layers. Both encoders have a linear projector $\mathrm{W_{p}}$ on top to map from each encoder’s representation to a shared space for contrastive learning.
@@
The label-wise image-concept matching (IWIC) score measures how the test image $\mathbf{x}$ matches a concept $\mathbf{y}_{c}$ and can be calculated using either global or local image features. The global IWIC score $s^{G}_{c}\left(\cdot\right)$ is defined as the cosine similarity between global image feature $h^{v}$ and the concept feature $h^{t}_{c}$: $s^{G}_{c}\left(\mathbf{x}\right)=\mathrm{cos\_sim}(h^{v},h^{t}_{c})$. The local IWIC score $s^{L}_{c}(\cdot)$ is the max-pooled cosine similarity of image patch features $p_{i}^{v}$ and the concept feature $h^{t}_{c}$: $s^{L}_{c}(\mathbf{x})=\max_{i}\mathrm{cos\_sim}(p_{i}^{v},h^{t}_{c})$. The MCM and GL-MCM scores are then defined as:
@@
where $\tau$ and $\tau^{\prime}$ are the temperature hyperparameters. MCM only uses global image feature, while GL-MCM additionally considers local image features. For ID data, both MCM and GL-MCM scores will be matched to one of the concept features with a high score; and vice versa. As a result, our OOD detection function can be formulated as:
@@ 3.1 OOD Detection with Selective Low-Rank Approximation
Motivated by the effectiveness of SVD in capturing the most informative aspects of a matrix, we propose a method to enhance OOD detection. By decomposing a weight matrix $W$ into its singular values and vectors, we can identify and retain the principle singular components that significantly contribute to the model’s performance. This approach ensures that the essential features of $W$ are preserved while discarding the less critical minor singular components. Given a weight matrix $W$ in CLIP (e.g., $\mathrm{W_{\text{up}}}$ or $\mathrm{W_{k}}$), we replace the matrix with its low-rank approximation part $\widehat{W}$ as described in Equation 5 (see Figure 1). Given the rank reduction ratio $\Theta$, the rank of $\widehat{W}$ is determined by $r(\widehat{W})=\text{round}((1-\Theta)\cdot r(W))$. This selective low-rank approximation leverages the compact representation provided by SVD to enhance the model’s ability to detect OOD instances effectively without requiring additional training. We demonstrate our method’s ability to improve OOD detection (Table 1) while maintaining ID classification performance (Table 7) in Section 4.2 and Section 4.5 .
@@
HyperParameter Search Algorithm
Due to the presence of many weight matrices in CLIP, each consisting of hundreds of singular values, conducting a complete search over all combinations of low-rank approximation weight matrices is impractical. Therefore, we propose a greedy search algorithm to determine the rank reduction ratio for each weight matrix. Among all linear weight matrices in each encoder layer, we focus on $\mathrm{W_{\text{up}}}$ as it is most effective according to our preliminary experiment. For simplicity, we assume both image and text encoders have $N$ encoder layers. As shown in Algorithm 1, we search by first enumerating all $N$ vision encoder layers sequentially from top to bottom and then all $N$ text encoder layers in the same way. This search order is concisely denoted as searching from $2N$ to the first layer in CLIP. We compare different search algorithms in Section 4.4. The rank reduction ratio for each layer is the objective in SeTAR which is searched among the candidate list $\mathbf{\Theta}=\{\Theta_{0},\Theta_{1},\cdots,\Theta_{J}\}$ according to the loss on the validation set. We employ the LoCoOp loss (Equation 12) proposed in (Miyai et al., 2023a) as our loss function. This loss requires only ID images. It contains an ID loss for ID image classification and an OOD loss to push away pseudo OOD features from the ID class text embeddings where the pseudo OOD features are from ID-irrelevant nuisances (Equation 10) (e.g., backgrounds) in CLIP’s local features. We refer the readers to Miyai et al. (2023a) or Appendix B for more details. For $\Theta_{j}\in\mathbf{\Theta}$, we remove $\Theta_{j}$ (in percent) singular values along with their corresponding singular vectors to obtain the approximated matrix $\mathrm{\widehat{W}_{\text{up}}}$ (Equation 5). It is worth noting that the rank reduction raio candidate list includes $\Theta_{0}=0$, indicating that the weight matrix has the chance to remain unmodified.
@@ 4.1 Experimental Setup
Following existing studies (Ming et al., 2022; Miyai et al., 2023b, a), we use CLIP ViT-B/16222https://huggingface.co/openai/clip-vit-base-patch16 (Radford et al., 2021) as our backbone. Both image and text encoders have 12 layers. More results with different backbones are in Section 4.4. The rank reduction ratio candidates range from 0 to 40% in 5% intervals. We use a temperature of 1111333Temperature is set to $1.0$ for the scaled CLIP logits, equivalent to the unscaled CLIP logits with a temperature of $100$. We adopt the unscaled setting in our implementation., unless stated otherwise. In all experiments, we use one CLIP text prompt: "a photo of a [CLASS],", where [CLASS] is the ID class name. We set hyperparameters $\lambda$ (Equation 12) and top-K (Equation 10) according to the specific ID datasets and backbones. Detailed settings are in Table 10, with a sensitivity analysis in Section 4.4. For SeTAR+FT and LoRA experiments, the learning rate and epoch number are set to $21e-$ 2 and $5$ for all experiments. The LoRA rank $r$ is set to is set to match e trainable parameters of SeTAR+FT. Detailed settings are in Table 11. We report results from three runs with seeds 3, 4, 5444For SeTAR , the results are the same under different random seeds as it does not require training.. All experiments are conducted on a single NVIDIA RTX 4090 GPU. The time cost for low-rank approximation with CLIP-base on the ImageNet1K validation set is about 20 minutes.
@@ 4.2 Training-free Results
The training-free OOD detection performances are summarized in Table 1. Compared with zero-shot baselines, a salient observation is that on both MCM and GL-MCM, using SeTAR outperforms the vanilla method by a large margin across all OOD detection tasks. For example, using Pascal-VOC as ID, SeTAR yields an average reduction of 12.84% FPR95 on MCM and 18.95% FPR95 on GL-MCM. Considering that SeTAR is generally applicable and training-free, these results are very encouraging. Comparing SeTAR with scoring function MCM and GL-MCM, SeTAR+GL-MCM performs better on all OOD detection tasks. However, the superiority of GL-MCM score over MCM appears to be contingent upon the choice of the model backbone. As evidenced in Table 8, SeTAR+MCM demonstrates superior performance with an average FPR95 reduction of 8.30% compared to SeTAR+GL-MCM with CLIP-large as the backbone on ImageNet1K.
@@ 4.3 Fine-tuning Results
In this section, we compare SeTAR+FT with fine-tuning baselines, including NPOS (Tao et al., 2023), CoOp (Zhou et al., 2022), LoCoOp (Miyai et al., 2023a) and LoRA (Hu et al., 2022). LoCoOp is the state-of-the-art prompt-learning OOD detection method on CLIP. LoRA is a representative parameter-efficient fine-tuning method. Followinging previous work (Tao et al., 2023; Zhou et al., 2022; Miyai et al., 2023a), we report the results on the the ImageNet1K benchmark in Table 2. We observe that SeTAR+FT outperforms all baselines on both MCM and GL-MCM scoring functions. For example, with CLIP-base as the backbone, SeTAR+FT achieves an average FPR95 reduction of 3.97% and 6.67% compared to LoCoOp and LoRA.
@@
Moreover, when scaled up to CLIP-large, SeTAR+FT outperforms LoCoOp and LoRA by 17.92% and 12.45% FPR95 on the same benchmark. Similar results are observed on Swin Transformer (Liu et al., 2021), where SeTAR+FT outperforms LoRA by 17.36% FPR95 on the MSP scoring function and 36.80% FPR95 on the Energy scoring function. These results demonstrate the effectiveness and scalability of SeTAR+FT in improving the OOD detection performance of CLIP.
Furthermore, we found that SeTAR+FT is more effective than LoRA, as demonstrated in Figure 6. The results show that SeTAR+FT achieves a quicker convergence and lower loss than LoRA during the fine-tuning process, especially OOD loss, indicating that SeTAR+FT is more effective in adapting the pre-trained weights to the OOD detection task.
@@ 4.4 Ablation Study
Table 3 shows an ablation study on the modality involved in SeTAR. As shown, the vision modality outperforms the text modality. When considering the vision modality alone and the combined vision+text modality, the latter either outperforms or achieves comparable average results to the former. Consequently, we make modifications to both the vision and text modalities in SeTAR to enhance overall performance.
@@
In this part, we present empirical evidence for modifying $\mathrm{W_{\text{up}}}$. We first compare the performance of SeTAR with different types of weight matrix in each Transformer layer, including $\mathrm{W_{q}}$, $\mathrm{W_{k}}$, $\mathrm{W_{v}}$, $\mathrm{W_{o}}$, $\mathrm{W_{\text{up}}}$ and $\mathrm{W_{\text{down}}}$. As shown in Figure 2 and Figure 3 of Appendx F, the $X$-axis denotes the number of weight matrixes (layers) that we have searched, while the $Y$-axis is the average AUROC and FPR95. The results show that $\mathrm{W_{\text{up}}}$ consistently outperforms other weight matrices in terms of both AUROC and FPR95.
@@
In addition to weight matrics in each transformer layer, CLIP has one projection matrix $\mathrm{W_{p}}$ on top of each encoder, which serves to project image/text representations into a shared space. In Table 4, we compare the performance of SeTAR with and without modifying $\mathrm{W_{p}}$. We search $\mathrm{W_{p}}$ first right before searching the image/text encoder. The results show that frozen $\mathrm{W_{p}}$ brings a reduction of $4.20\%$ FPR95. Consequently, we keep $\mathrm{W_{p}}$ frozen in SeTAR.
@@
Inspired from SVD, SeTAR modified the model weights by pruning the minor singular components, and retains the principle components that contribute the most to the model’s performance. To validate this design, we compare SeTAR with two alternatives: principal component pruning and random pruning. Principal component pruning takes
@@
an opposite approach, retaining minor components and pruning major ones. Random pruning, on the other hand, prunes weights randomly. As shown in Table 6, principle pruning suffers from a significant performance drop compared to SeTAR , while random pruning performs slightly better than principle pruning. These results demonstrate the effectiveness of SeTAR ’s design choice in pruning the minor components.
@@
In this section, we present the sensitivity analysis of the hyperparameters $\lambda$ (Figure 4) and top-K (Figure 5). As observed in Figure 4, the average AUROC remains stable at lower values and slightly decreases as $\lambda$ increases for both SeTAR+MCM and SeTAR+GL-MCM. Notably, the optimal setting of $\lambda$ may vary depending on the model backbone, with our experiments indicating that CLIP-large may require a larger $\lambda$ than CLIP-base. Regarding top-K, which controls the number of OOD regions considered, higher top-K values include more OOD regions. When top-K equals the number of ID classes, the model considers all OOD regions. Conversely, when top-K is set to 0, the model focuses solely on ID loss. Figure 5 demonstrates the impact of varying top-K on AUROC and FPR95. SeTAR exhibits high robustness to changes in top-K, except at extreme values such as 0 or the maximum number of ID classes.

### v2

@@ Abstract
Out-of-distribution (OOD) detection is crucial for the safe deployment of neural networks. Existing CLIP-based approaches perform OOD detection by devising novel scoring functions or sophisticated fine-tuning methods. In this work, we propose SeTAR, a novel, training-free OOD detection method that leverages selective low-rank approximation of weight matrices in vision-language and vision-only models. SeTAR enhances OOD detection via post-hoc modification of the model’s weight matrices using a simple greedy search algorithm. Based on SeTAR, we further propose SeTAR+FT, a fine-tuning extension optimizing model performance for OOD detection tasks. Extensive evaluations on ImageNet1K and Pascal-VOC benchmarks show SeTAR’s superior performance, reducing the relatively false positive rate by up to 18.95% and 36.80% compared to zero-shot and fine-tuning baselines. Ablation studies further validate SeTAR’s effectiveness, robustness, and generalizability across different model backbones. Our work offers a scalable, efficient solution for OOD detection, setting a new state-of-the-art in this area.
@@ 1 Introduction
Previous research has primarily focused on detecting OOD instances in either visual (DeVries & Taylor, 2018; Liang et al., 2018; Hendrycks et al., 2022) or textual data (Hu & Khan, 2021; Zheng et al., 2020; Zhou et al., 2021). Recently, significant progress has been made in multimodal tasks like multimodal retrieval (Li et al., 2023; Caesar et al., 2018) and image classification (Yu et al., 2022), thanks to vision-and-language pretrained (VLP) models like CLIP (Radford et al., 2021). More recent studies have explored OOD detection with CLIP, grouped into zero-shot methods (Fort et al., 2021; Ming et al., 2022; Miyai et al., 2023b) and finetuning-based methods (Ming & Li, 2023; Tao et al., 2023; Miyai et al., 2023a). However, the zero-shot methods suffer from suboptimal performance due to potential domain gaps with ID downstream data. On the other hand, finetuning-based methods carry the risk of deconstructing the intricate representations learned by CLIP which requires a meticulously designed training strategy. Sparsification-based approaches (Sun et al., 2021; Djurisic et al., 2023) have demonstrated potential in OOD detection within CNNs, leveraging the assumption that ID and OOD samples produce distinct activation patterns. Nevertheless, their effectiveness diminishes in large-scale pre-trained models such as CLIP, where activation differences become more subtle, thereby limiting their applicability primarily to models fine-tuned on downstream ID-domain datasets.
@@
We conduct extensive evaluations and achieve state-of-the-art performance on common OOD detection benchmarks for CLIP, including the ImageNet1K and Pascal-VOC benchmarks. Compared to vanilla MCM and GL-MCM, SeTAR with the CLIP backbone reduces relatively FPR95 by 9.5% and 12.0% on average across two benchmarks, respectively. When further integrate fine-tuning into SeTAR, SeTAR+FT outperforms the state-of-the-art fine-tuning baselines LoCoOp (Miyai et al., 2023a) and LoRA (Hu et al., 2022). Moreover, we perform a comprehensive ablation study and analysis to verify and understand SeTAR. In summary, our key results and contributions:
@@
    We proposed SeTAR, a simple yet effective OOD detection method based on selective low-rank approximation. It is training-free as it only performs post-hoc modification to weight matrices. SeTAR applies to a variety of scoring functions and model backbones. It can be readily integrated with existing zero-shot OOD detection methods.
@@ 2 Preliminaries
The CLIP model (Radford et al., 2021) comprises an image encoder $E^{v}(\cdot)$ and a text encoder $E^{t}(\cdot)$, aligned via contrastive learning on web-scale image-text pairs. We focus on CLIP-ViT, where the image encoder is a Vision Transformer (ViT). Each ViT layer includes a multihead self-attention sublayer and a feed-forward sublayer. In the self-attention module, the hidden state is projected into different spaces using learnable parameters $\mathrm{W_{q},W_{k},W_{v}}$. The outputs are concatenated and projected back with another linear matrix $\mathrm{W_{o}}$. The feed-forward module projects the hidden state into a wider space using $\mathrm{W_{\text{up}}}$ and then back with $\mathrm{W_{\text{down}}}$ after a non-linear activation (Figure 1). Given the similarity between the image and text encoder layers, we adopt consistent notations for the linear matrices in both. Each encoder also includes a linear projector $\mathrm{W_{p}}$ to map their representations into a shared space for contrastive learning.
@@
The label-wise image-concept matching (IWIC) score measures how well a test image $\mathbf{x}$ aligns with a concept $\mathbf{y}_{c}$, using either global or local features. The global IWIC score $s^{G}_{c}\left(\cdot\right)$ is the cosine similarity between the global image feature $h^{v}$ and the concept feature $h^{t}_{c}$: $s^{G}_{c}\left(\mathbf{x}\right)=\mathrm{cos\_sim}(h^{v},h^{t}_{c})$. The local IWIC score $s^{L}_{c}(\cdot)$ is the max-pooled cosine similarity between image patch features $p_{i}^{v}$ and the concept feature $h^{t}_{c}$: $s^{L}_{c}(\mathbf{x})=\max_{i}\mathrm{cos\_sim}(p_{i}^{v},h^{t}_{c})$. The MCM and GL-MCM scores are defined as:
@@
where $\tau$ and $\tau^{\prime}$ are the temperature hyperparameters. MCM only uses global image features, while GL-MCM additionally considers local image features. For ID data, both MCM and GL-MCM scores will be matched to one of the concept features with a high score; and vice versa. As a result, our OOD detection function can be formulated as:
@@ 3.1 OOD Detection with Selective Low-Rank Approximation
SVD-based weight pruning, particularly in noise-prone layers, can substantially reduce a network’s sensitivity to minor perturbations, leading to enhanced stability and robustness (Yao et al., 2024). This stability is crucial for OOD detection, as it ensures the model’s reliable performance across a wide range of inputs. Building on this, we propose a method to improve OOD detection by selectively applying low-rank approximation to weight matrices. By decomposing a weight matrix $W$ into its singular values and vectors, we can identify and retain the principle singular components that significantly contribute to the model’s performance. This approach ensures that the essential features of $W$ are preserved while discarding the less critical minor singular components. Given a weight matrix $W$ in CLIP (e.g., $\mathrm{W_{\text{up}}}$ or $\mathrm{W_{k}}$), we replace the matrix with its low-rank approximation part $\widehat{W}$ as described in Equation 5 (see Figure 1). Given the rank reduction ratio $\Theta$, the rank of $\widehat{W}$ is determined by $r(\widehat{W})=\text{round}((1-\Theta)\cdot r(W))$. This selective low-rank approximation leverages the compact representation provided by SVD to enhance the model’s ability to detect OOD instances effectively without requiring additional training. We demonstrate our method’s ability to improve OOD detection (Table 1) while maintaining ID classification performance (Table 7) in Section 4.2 and Section 4.5.
HyperParameter Search Algorithm
Due to the presence of many weight matrices in CLIP, each consisting of hundreds of singular values, conducting a complete search over all combinations of low-rank approximation weight matrices is impractical. Therefore, we propose a greedy search algorithm to determine the rank reduction ratio for each weight matrix. Among all linear weight matrices in each encoder layer, we focus on $\mathrm{W_{\text{up}}}$ as it is most effective according to our preliminary experiment. For simplicity, we assume both image and text encoders have $N$ encoder layers. As shown in Algorithm 1, we search by first enumerating all $N$ vision encoder layers sequentially from top to bottom and then all $N$ text encoder layers in the same way. This search order is concisely denoted as searching from $2N$ to the first layer in CLIP. We compare different search algorithms in Section 4.4. The rank reduction ratio for each layer is the objective in SeTAR which is searched among the candidate list $\mathbf{\Theta}=\{\Theta_{0},\Theta_{1},\cdots,\Theta_{J}\}$ according to the loss on the validation set. We employ the LoCoOp loss (Equation 12) proposed in (Miyai et al., 2023a) as our loss function. This loss requires only ID images. It contains an ID loss for ID image classification and an OOD loss to push away pseudo OOD features from the ID class text embeddings where the pseudo OOD features are from ID-irrelevant nuisances (Equation 10) (e.g., backgrounds) in CLIP’s local features. We refer the readers to Miyai et al. (2023a) or Appendix B for more details.
@@
For $\Theta_{j}\in\mathbf{\Theta}$, we remove $\Theta_{j}$ (in percent) singular values along with their corresponding singular vectors to obtain the approximated matrix $\mathrm{\widehat{W}_{\text{up}}}$ (Equation 5). It is worth noting that the rank reduction raio candidate list includes $\Theta_{0}=0$, indicating that the weight matrix has the chance to remain unmodified.
@@ 4.1 Experimental Setup
Following existing studies (Ming et al., 2022; Miyai et al., 2023b, a), we use CLIP ViT-B/16222https://huggingface.co/openai/clip-vit-base-patch16 (Radford et al., 2021) as our backbone. Both image and text encoders have 12 layers. More results with different backbones are in Section 4.4. The rank reduction ratio candidates range from 0 to 40% in 5% intervals. We use a temperature of 1111333Temperature is set to $1.0$ for the scaled CLIP logits, equivalent to the unscaled CLIP logits with a temperature of $100$. We adopt the unscaled setting in our implementation., unless stated otherwise. In all experiments, we use one CLIP text prompt: "a photo of a [CLASS],", where [CLASS] is the ID class name. We set hyperparameters $\lambda$ (Equation 12) and top-K (Equation 10) according to the specific ID datasets and backbones. Detailed settings are in Table 12, with a sensitivity analysis in Section 4.4. For SeTAR+FT and LoRA experiments, the learning rate and epoch number are set to $21e-$ 2 and $5$ for all experiments. The LoRA rank $r$ is set to match the trainable parameters of SeTAR+FT. Detailed settings are in Table 13. We report results from three runs with seeds 3, 4, 5444For SeTAR , the results are the same under different random seeds as it does not require training.. All experiments are conducted on a single NVIDIA RTX 4090 GPU. The time cost for low-rank approximation with CLIP-base on the ImageNet1K validation set is about 20 minutes.
@@ 4.2 Training-free Results
The training-free OOD detection performances are summarized in Table 1. Compared with zero-shot baselines, a salient observation is that on both MCM and GL-MCM, using SeTAR outperforms the vanilla method by a large margin across all OOD detection tasks. For example, using Pascal-VOC as ID, SeTAR yields a relatively average reduction of 12.84% FPR95 on MCM and 18.95% FPR95 on GL-MCM. Considering that SeTAR is generally applicable and training-free, these results are very encouraging. Comparing SeTAR with scoring function MCM and GL-MCM, SeTAR+GL-MCM performs better on all OOD detection tasks. However, the superiority of GL-MCM score over MCM appears to be contingent upon the choice of the model backbone. As evidenced in Table 8, SeTAR+MCM demonstrates superior performance with a relatively average FPR95 reduction of 8.30% compared to SeTAR+GL-MCM with CLIP-large as the backbone on ImageNet1K.
@@ 4.3 Fine-tuning Results
In this section, we compare SeTAR+FT with fine-tuning baselines, including NPOS (Tao et al., 2023), CoOp (Zhou et al., 2022), LoCoOp (Miyai et al., 2023a) and LoRA (Hu et al., 2022). LoCoOp is the state-of-the-art prompt-learning OOD detection method on CLIP. LoRA is a representative parameter-efficient fine-tuning method. Following previous work (Tao et al., 2023; Zhou et al., 2022; Miyai et al., 2023a), we report the results on the the ImageNet1K benchmark in Table 2. We observe that SeTAR+FT outperforms all baselines on both MCM and GL-MCM scoring functions. For example, with CLIP-base as the backbone, SeTAR+FT achieves a relatively average FPR95 reduction of 3.97% and 6.67% compared to LoCoOp and LoRA. Moreover, when scaled up to CLIP-large, SeTAR+FT outperforms LoCoOp and LoRA by
@@
relatively 17.92% and 12.45% FPR95 on the same benchmark. Similar results are observed on Swin Transformer (Liu et al., 2021), where SeTAR+FT outperforms LoRA by relatively 17.36% and 36.80% FPR95 on MSP and Energy scoring functions, respectively. The larger improvement on Swin Transformer may stem from its reliance on ImageNet training, making it prone to overfitting and weaker at OOD detection. Our method mitigates these issues, enhancing Swin’s generalization to OOD instances. These results demonstrate the effectiveness and scalability of SeTAR+FT in improving the OOD detection performance.
Furthermore, as shown in Figure 6, SeTAR+FT demonstrates faster convergence and lower loss than LoRA, especially in OOD loss, indicating that SeTAR+FT is more effective in adapting the pre-trained weights to the OOD detection task.
@@ 4.4 Ablation Study
Table 3 shows an ablation study on the modality involved in SeTAR.
@@
As shown, the vision modality outperforms the text modality, indicating the vision modality is more dominant in enhancing the model’s performance. When considering the vision modality alone and the combined vision+text modality, the latter either outperforms or achieves comparable average results to the former. Consequently, we make modifications to both the vision and text modalities in SeTAR to enhance overall performance.
@@
In this part, we present empirical evidence for modifying $\mathrm{W_{\text{up}}}$. We first compare the performance of SeTAR with different types of weight matrix in each Transformer layer, including $\mathrm{W_{q}}$, $\mathrm{W_{k}}$, $\mathrm{W_{v}}$, $\mathrm{W_{o}}$, $\mathrm{W_{\text{up}}}$ and $\mathrm{W_{\text{down}}}$. As shown in Figure 2 and Figure 3 of Appendx F, the $X$-axis denotes the number of weight matrixes (layers) that we have searched, while the $Y$-axis is the average AUROC and FPR95. The results show that $\mathrm{W_{\text{up}}}$ consistently outperforms other weight matrices in terms of both AUROC and FPR95. In addition to weight matrics in each transformer layer, CLIP has one projection matrix $\mathrm{W_{p}}$ on top of each encoder, which serves to project image/text representations into a shared space. In Table 4, we compare the performance of SeTAR with and without modifying $\mathrm{W_{p}}$. We search $\mathrm{W_{p}}$ first right before searching the image/text encoder. The results show that frozen $\mathrm{W_{p}}$ brings a relatively reduction of $4.20\%$ FPR95. Consequently, we keep $\mathrm{W_{p}}$ frozen in SeTAR.
@@
Inspired from SVD, SeTAR modified the model weights by pruning the minor singular components, and retains the principle components that contribute the most to the model’s performance. To validate this design, we compare SeTAR with two alternatives: principal component pruning and random pruning pruning. Principal component takes
@@
the opposite approach, retaining minor components and pruning major ones. Random pruning, on the other hand, prunes weights randomly. As shown in Table 6, principle pruning suffers from a significant performance drop compared to SeTAR , while random pruning performs slightly better than principle pruning. These results demonstrate the effectiveness of SeTAR ’s design choice in pruning the minor components.
@@
In this section, we present the sensitivity analysis of the hyperparameters $\lambda$ (Figure 4) and top-K (Figure 5). As observed in Figure 4, the average AUROC remains stable at lower values and slightly decreases as $\lambda$ increases for both SeTAR+MCM and SeTAR+GL-MCM. Notably, the optimal setting of $\lambda$ may vary depending on the model backbone, with our experiments indicating that CLIP-large may require a larger $\lambda$ than CLIP-base. Despite this variation, the $\lambda$ parameter demonstrates strong transferability across datasets for the same backbone. Swapping the optimal $\lambda$ between ImageNet1K and Pascal-VOC has a minimal performance impact, consistently outperforming the vanilla method. With the VOC-optimized $\lambda$ on ImageNet1K, CLIP-base achieves an FPR95 of 40.91 and AUROC of 91.02, and CLIP-large reaches 46.73 FPR95 and 91.81 AUROC. Conversely, using the ImageNet1K-optimized $\lambda$ on Pascal-VOC, CLIP-base achieves 33.18 FPR95 and 93.65 AUROC, while CLIP-large attains 44.39 FPR95 and 92.3 AUROC.
Top-K controls the number of OOD regions considered: higher values include more OOD regions, with top-K equal to the number of ID classes covering all OOD regions, and top-K set to 0 focusing solely on ID loss. The optimal top-K depends on the number of ID categories, making it non-transferable across datasets. However, SeTAR remains robust to top-K variations, as shown in Figure 5, except at extreme values (0 or the maximum number of classes). We recommend setting top-K to around 30% of the total categories, such as 300 for ImageNet1K and 4 for Pascal-VOC. For the Swin-base model, top-K at 300 on ImageNet1K yields an FPR95 of 56.82 and AUROC of 85.68 with MSP, and an FPR95 of 52.56 and AUROC of 84.51 with Energy.